On the Zero modes of Pauli and Dirac operators

W. Desmond Evans Cardiff University, Wales, UK

The talk will present some recent results obtained jointly with A. Balinsky which prove that, within an appropriate class, magnetic fields which give rise to zero modes are rare. An estimate of the number of zero modes is also given.